# Nevanlinna Theory of Algebroid Functions on Complete Kähler Manifolds

### v9

@@ Abstract.
We generalize the classical Nevanlinna theory for algebroid functions on to a complete Kähler manifold with non-negative Ricci or non-positive sectional curvature. As an application, we give some Picard type theorems and five-value type theorems for algebroid functions under certain conditions.
@@ 1.1. Motivation
The notion of algebroid functions (see [16, 28]) introduced by H. Poincaré is a generalization of algebraic functions and meromorphic functions. Let us review that a -valued algebroid function on a (nonempty) domain (the complex plane) is defined by an irreducible algebraic equation
where are holomorphic functions with no common zeros on An algebroid function is viewed as a special multi-valued analytic function, while which was first put forward, G. Darboux deemed that it is an important class of functions. To see that, for example, it is known that a -valued algebroid function defines a -leaf Riemann surface which is a (generic) -sheeted analytic covering of A further fact claims that can lift to a single-valued meromorphic function on via a natural projection such that which is called the uniformization of
The most concerned problem on algebroid functions seems to be the value distribution. In the 1930s after R. Nevanlinna [22] founded two fundamental theorems concerning the value distribution theory of meromorphic functions in 1925, many famous scholars such as G. Rémoundos, H. Selberg, E. Ullrich and G. Valiron, etc. began to study the value distribution theory of algebroid functions. For example, G. Rémoundos [24] proved a Picard’s theorem which claims that a nonconstant -valued algebroid function on omits distinct values at most in Afterwards, G. Valiron [40] showed without details a five-value type theorem in unicity theory, which claims that two nonconstant -valued algebroid functions on will be identical if they share +1 distinct values ignoring multiplicities in Since then, a large number of scholars have devoted themselves to this topic, including investigations to Nevanlinna theory, Ahlfors theory of covering surfaces, unicity theory and normal family theory, etc. More details refer to H. Cartan [7], He-Xiao [16], Niino-Ozawa [19], J. Noguchi [20], M. Ozawa [23], M. Ru [25, 26], Sun-Gao [28], J. Suzuki [29], L. Selberg [30, 31], N. Toda [37, 38] and E. Ullrich [39], etc.
Algebroid function theory in one complex variable didn’t mature until the last 30-40 years. However, there is little literature on the algebroid functions in several complex variables, except for the work of Hu-Yang [15]. Given two nonconstant algebroid functions on a non-compact Hermitian (Kähler) manifold . We are inspired by the following three fundamental problems:
@@
In this paper, we answer these three problems by generalizing the classical Nevanlinna theory for algebroid functions on to a non-compact complete Kähler manifold with non-negative Ricci or non-positive sectional curvature.
@@ A. Uniformization
Let be a non-compact Hermitian manifold of complex dimension At first, we generalize the notion of algebroid functions from to by means of a technical treatment (see Definition 2.15). Let be a -valued algebroid function on Assume that is a -leaf meromorphic function element of namely, is a local -valued meromorphic solution of equation (7) in a neighborhood of where is a branching point with order of If is a “non-singular” branching point of then we describe the asymptotic behavior of as approaches i.e., we prove that can be expanded into the Pusieux series (see Theorem 2.14)
with around in local holomorphic coordinates where are meromorphic functions in With this Pusieux series, we show that
@@
Let be the -leaf complex manifold of a -valued algebroid function on Then, can lift to a meromorphic function on via the natural projection such that
@@ B. Nevanlinna theory
In this part and the following Parts C, D, we utilize to stand for a non-compact complete Kähler manifold. Let be a -valued algebroid function on Refer to the definition in Section 5.3, we have the Nevanlinna’s functions of and on where is the Ricci form of and is a precompact domain containing a fixed reference point To unify the form, we write
@@
Theorem B4 covers the classical second main theorem for both and the unit ball with Poincaré metric in (see Corollaries 5.7 and 5.8).
@@ C. Propagation of algebraic dependence
Let be the hyperplane line over We denote by the set of all indecomposable hypersurfaces in such that for some where and is the natural projection on the -th factor. Put
@@ 2.1. Notion of Algebroid Functions
First of all, we would like to extend the notion of algebroid functions to a non-compact Hermitian manifold. Let be a -dimensional non-compact Hermitian manifold. Consider an irreducible algebraic equation in on :
@@
Let be the indeterminacy locus of which is an analytic set in of complex codimension not less than if or an empty set if So, is a holomorphic mapping of into
@@
where is an open covering of Whence, we see that is well defined on and it hence defines a meromorphic function on for
@@ 2.1.1. Meromorphic function elements
Let A pair is called a regular function element, if is holomorphic in a neighborhood of A pair is called a meromorphic function element, if is meromorphic in a neighborhood of A pair is said to be subordinate to equation if is a local solution of equation in a neighborhood of
@@
Then there exists a unique regular function element subordinate to equation satisfying
@@
The theorem can be showed by employing implicit function theorem. Here, we present an alternative proof of complex analysis. Taking a such that (see Remark 2.1). We write in the form
@@
The conditions mean that
By the continuation, there exist such that
for all satisfying and In addition, those are in In further, restrict to then we obtain
Employing Rouché’s theorem, for any fixed with and have the same number of zeros in the disc In fact, the number of zeros is 1, since in has only one zero in the disc when This implies that uniquely determines a function in a smaller geodesic ball neighborhood of with radius such that
Evidently, is continuous on
Next, we show that is holomorphic on One just needs to show without loss of generality that is analytic at Taking local holomorphic coordinates around such that Apply Hartog’s theorem, it is sufficient to show that is holomorphic in every holomorphic direction at For and , set
@@
has exactly distinct complex roots. Therefore, Theorem 2.3 implies that
@@
where for According to Sylvester’s elimination method, we can see that if and only if or has a multiple root as a polynomial in i.e., This gives that is an analytic set of complex dimension Define the discriminant of by
@@
Since then one can see that is a zero of i.e., a pole of In fact, we will deduce from Theorem 2.22 (in Section 2.3) that if and only if To conclude, we see that Corollary 2.5 can be extended to meromorphic function elements as follows:
@@ 2.1.2. Analytic continuation of meromorphic function elements
Recall that a meromorphic function element on is a pair where is a meromorphic function defined in a neighborhood of in As a matter of convenience, we write
for short without confusion. Instead of we sometimes take a geodesic ball neighborhood centered at with radius Let be any two meromorphic function elements on Define an equivalent relation as follows: if and only if
@@
We equip with a topology. A meromorphic function element is said to be a direct analytic continuation of a meromorphic function element if and in a neighborhood of By the analytic continuation, we can regard that
A neighborhood of is defined by a set (containing ) of all such that is a direct analytic continuation of satisfying where is a neighborhood of in Namely, In particular, a -neighborhood of (satisfying that ), is defined by the set (which contains ) of all such that is a direct analytic continuation of satisfying that In other words, it is defined to be A subset is called an open set, if either is an empty set or every point is an inner point, i.e., there exists a -neighborhood of such that It is not hard to check that it defines a topology of
@@
Pick If then we can consider -neighborhoods of and of where Evidently, we have If for then we can take a function element It implies that in a small neighborhood of By the uniqueness theorem of analytic functions, we obtain but it contradicts with ∎
Next, we introduce the notion of analytic continuation (by which we mean both holomorphic continuation and meromorphic continuation) of meromorphic function elements.
@@
The analytic continuation of elements in along a curve in is a continuous mapping of into namely, an analytic continuation of to along is a curve connecting and Moreover, if is an analytic continuation of along a curve, then is also an analytic continuation of along the same curve. Thus, the analytic continuation is symmetric.
@@
Then, we have We show Otherwise, we assume that By Corollary 2.6, there are exactly function elements at say that
Take small enough such that where Then there is a unique integer such that Notice that the domain of definition of is so for any one can redefine function elements Combined with the definition of it yields immediately that can continue analytically to along i.e., is an analytic continuation of along However, it contradicts with the definition of Thus, we obtain The uniqueness of is immediate due to the uniqueness theorem of analytic (meromorphic) functions. ∎
@@
Let be distinct function elements at respectively. Then can continue analytically to each other along any curve in
@@
Let be distinct function elements at every fixed With the aid of Corollary 2.11, then we only need to prove that all can continue analytically to each other. Let be the largest integer such that there are function elements in which can continue analytically to each other along a curve. Clearly, we have Without loss of generality, we shall assume that these function elements are So, we can write
where are holomorphic functions in a neighborhood of By Vieta’s theorem, we obtain
It is equivalent to show that: can continue analytically to the whole Apply Theorem 2.10, can continue analytically to along a curve at any where is a permutation of By Vieta’s theorem, determine the same in a small neighborhood of i.e., continue analytically to Hence, holds. If then we are able to determine another set via the rest function elements However, it implies that equation (7) is reducible, which is a contradiction. ∎
@@
2.1.3. Properties of meromorphic function elements near multiple points
We discuss the local properties of function elements near multiple points. Let be a pair satisfying that and
@@ 2.1.3. Properties of meromorphic function elements near multiple points
Taking a sufficiently small neighborhood of such that on By the continuation of there exists a number such that on Let then there exist distinct regular function elements and in respectively. By the use of Theorem 2.10, then for a curve with and are able to continue analytically to along In other words, there exists a permutation of such that can continue analytically to along respectively. By symmetrical property, can also continue analytically to along However, will not, in general, continue analytically back to along respectively. In the following, we discuss the analytic continuation of a function element along a closed Jordan curve surrounding a multiple point. Take Pick as a closed Jordan curve passing by around Without loss of generality, we only extend analytically along starting from When returns to from along one will attain to a function If then one continues to extend analytically along Repeating this action, we see finally that cycles on and on periodically. Let be the smallest period such that returns to itself, we obtain We say that is a branching point with order (multiplicity) of if Repeatedly, then it is not hard to verify that can be divided into several groups. Let be the number of groups. If the -th group has members, then we have
Without loss of generality, we consider the first group: Take a proper we have a biholomorphic mapping such that
@@
where stands for the polydisc centered at with polyradius in and is a local holomorphic coordinate of in So, we have
Set and write Next, we shall show that is a single-valued analytic function on where is the unit disc centered at in the -plane. Evidently, is holomorphic on We still need to prove that is a removed singularity of In doing so, it is sufficient to show that is bounded on due to Riemann’s analytic continuation. There exists a number such that on for Combining (7) with that on we obtain
@@
on If we obtain Thus, on and this bound does not depend on Whence, is bounded on This means that is holomorphic on and of the Taylor expansion
@@
near Let be the set of all branching points of in Note that the value of is independent of if is sufficiently small. If is a non-singular branching point of i.e., does not lie in a crossing of components of then is smooth whenever is small enough. By and Riemann’s analytic continuation (because the boundedness of on and is properly chosen if necessary), there is a biholomorphic mapping with
@@
such that can be expanded into the Pusieux series
near where are holomorphic functions in satisfying that for
@@
According to the above arguments, we show that is a -valued analytic function near In other words, if we cut along a (real) hypersurface which passes through then is separated into single-valued analytic components, and these components replace each other when they continue analytically along a closed curve around in By a local expansion of is a -multiple -valued branching point (with order ) of
Assume that is not a singular branching point of Set and write
Similarly, we have
If then
Choose such that Put then
Thus, i.e., the case turns to the case Note that is a pole with order of
@@
To conclude, for each we have confirmed that there are distinct (single-valued) meromorphic function elements subordinate to equation (7), if is not a branching point. If is a branching point, then one shows that there are several multi-valued function elements (which contain single-valued function elements) subordinate to equation (7), say where is a -order branching point of satisfying At this time, we call a -order branching point of If is a non-singular branching point of then we have the convergent Pusieux expansion
near in a local holomorphic coordinate Note that is a -multiple root of namely, can separate into single-valued components with value at If then can separate into single-valued components which take at In particular, is a meromorphic function element for Furthermore, is a regular function element if also. By the analytic continuation, equation (7) uniquely defines a -valued meromorphic function on outside a set of branching points of and is separated into single-valued meromorphic components on Let denote the set of all branching points of which is an analytic set contained in called the branching set of As a result, we conclude that
@@
Let be a -valued function element subordinate to equation If is not a singular branching point of then can be expanded into the convergent Pusieux series
in a local holomorphic coordinate near such that =0 and for where is the branching set in a local neighborhood of and are meromorphic functions in
We conclude the definition of algebroid functions on as follows.
@@
where are holomorphic functions locally defined on such that they well define a meromorphic mapping from into
Let denote distinct (single-valued) meromorphic components of Given any -leaf function element of with If denote the single-valued components of then there exists a permutation of such that on In other words, defines a -leaf function element on Thus, any function element of can be represented locally by one or several single-valued components of
2.2. Multi-leaf Complex Manifolds of Algebroid Functions
Let be a function element. A neighborhood of is defined as a set of all such that is a direct analytic continuation of satisfying i.e., Taking a proper if necessary, there exists a biholomorphic mapping such that where is the unit ball centered at the coordinate origin in Define a natural mapping
via
Evidently, is a homeomorphism. Its inverse is
Set We show that is a complex atlas of Let It suffices to show the transition mapping

### v10

@@ Abstract.
In this paper, we generalize the classical Nevanlinna theory of algebroid functions from to a complete Kähler manifold with either non-negative Ricci curvature or non-positive sectional curvature. As its applications, we establish some Picard type theorems and five-value type theorems for algebroid functions under certain conditions.
@@ 1.1. Motivation
The notion of algebroid functions (see [16, 28]) introduced by H. Poincaré is an extension of meromorphic functions. Recall that a -valued algebroid function on a domain is defined by an irreducible algebraic equation
where are holomorphic functions without common zeros in An algebroid function is viewed as a special multi-valued analytic function, while which was first put forward, G. Darboux deemed that it is an important class of functions. To see that, for example, it is known that a -valued algebroid function defines a -leaf Riemann surface which is a generic -sheeted analytic covering of A further fact states that can lift to a single-valued meromorphic function on via a natural projection such that which is called the uniformization of
The most concerned problem on algebroid functions seems to be the value distribution. In the 1930s after R. Nevanlinna [22] founded two fundamental theorems concerning the value distribution theory of meromorphic functions in 1925, many famous scholars such as G. Rémoundos, H. Selberg, E. Ullrich and G. Valiron, etc. began to study the value distribution theory of algebroid functions. For example, G. Rémoundos [24] proved a Picard’s theorem which claims that a nonconstant -valued algebroid function on omits distinct values at most in Afterwards, G. Valiron [40] showed without details a five-value type theorem in unicity theory, which claims that two nonconstant -valued algebroid functions on will be identical if they share +1 distinct values ignoring multiplicities in Since then, a large number of scholars have devoted themselves to the study of Nevanlinna theory, Ahlfors theory of covering surfaces, unicity theory and normal family theory, etc. for algebroid functions. We refer readers to H. Cartan [7], He-Xiao [16], Niino-Ozawa [19], J. Noguchi [20], M. Ozawa [23], M. Ru [25, 26], Sun-Gao [28], J. Suzuki [29], L. Selberg [30, 31], N. Toda [37, 38] and E. Ullrich [39], etc.
Algebroid function theory in one complex variable didn’t mature until the last 30-40 years. Up to now, there seems to be very little literature regarding algebroid functions in several complex variables, except for the limited work by Hu-Yang [15]. Let denote any two non-constant algebroid functions on a non-compact Hermitian or Kähler manifold We are inspired by the following three fundamental problems:
@@
In the present paper, we answer these problems. What is more important is that we develop the classical Nevanlinna theory of algebroid functions on to complete non-compact Kähler manifolds.
@@ A. Uniformization
Let be a non-compact Hermitian manifold. We generalize the notion of algebroid functions from to (see Definition 2.15). Let be a -valued algebroid function on Given a -leaf function element of When is not a singular branching point of we give a local Pusieux expansion of at i.e., in a local holomorphic coordinate we show that can be expanded locally into the following convergent Pusieux series (see Theorem 2.14)
at with where are meromorphic functions. By this Pusieux series, we construct a -leaf complex manifold and prove that
@@
Let be the -leaf complex manifold by a -valued algebroid function on Then, can lift to a meromorphic function on via the natural projection such that
@@ B. Nevanlinna theory
In this part and the following Parts C, D, we use to stand for a complete non-compact Kähler manifold. Let be any -valued algebroid function on Refer to the definition in Section 5.3, we have the Nevanlinna’s functions of and on where is the Ricci form of and is a precompact domain containing a fixed point To unify the form, we write
@@
Note that Theorem B4 covers the classical second main theorem for both and the unit ball with Poincaré metric in (see Corollaries 5.7 and 5.8).
@@ C. Propagation of algebraic dependence
Let be the point line bundle over We denote by the set of all indecomposable hypersurfaces of such that for some in which and is the natural projection on the -th factor. Set
@@ 2.1. Notion of Algebroid Functions
First of all, we would like to extend the notion of algebroid functions to a non-compact Hermitian manifold. Let be an -dimensional non-compact Hermitian manifold. Consider an irreducible algebraic equation in on :
@@
Let be the indeterminacy locus of which is either an analytic set of complex codimension not smaller than for or an empty set for in Then, is holomorphic on
@@
where is an open covering of Whence, we see that is well defined on and it hence defines a meromorphic function on with
@@ 2.1.1. Meromorphic function elements
Let be any point in A pair is called a regular function element, if is a holomorphic function in a neighborhood of and called a meromorphic function element, if is a meromorphic function in a neighborhood of A pair is said to be subordinate to equation if is a local solution of equation in a neighborhood of
@@
Then, there exists a unique regular function element subordinate to equation satisfying
@@
The theorem can be confirmed by applying implicit function theorem. Here, we give an alternative proof in complex analysis. Take a such that given by Remark 2.1. Write as the form
@@
By the conditions, we have
The continuity implies that here exist such that
for all with and In further, restrict to then we obtain
According to Rouché’s theorem, we deduce that for all with and have the same number of zeros in the disc In fact, this number of zeros is 1, since in has the unique zero in the disc for all with It implies that uniquely determines a function in a smaller geodesic ball neighborhood of with radius such that Evidently, is continuous on
Next, we show that is holomorphic on One just needs to show without loss of generality that is analytic at Taking a local holomorphic coordinate near such that Thanks to Hartog’s theorem, it is sufficient to examine that is holomorphic in each holomorphic direction at For and put
@@
has exactly distinct complex roots. Hence, Theorem 2.3 implies that
@@
where for Apply Sylvester’s elimination method, we can deduce that if and only if or has a multiple root as a polynomial in i.e., Thus, it means that is an analytic set of complex codimension Define the discriminant of by
@@
Since we see that is a zero of i.e., a pole of In fact, we note from Theorem 2.21 (see Section 2.3) that if and only if To conclude, Corollary 2.5 can be extended to as follows.
@@ 2.1.2. Analytic continuation of meromorphic function elements
We recall that a meromorphic function element on is a pair such that is a meromorphic function defined in a neighborhood of For convenience, we write without any confusion. We usually take which is a geodesic ball neighborhood centered at with radius For two meromorphic function elements and define an equivalent relation: if and only if
@@
We equip with a topology. A meromorphic function element is said to be a direct analytic continuation of a meromorphic function element if and in a neighborhood of By the analytic continuation, we may regard that
A neighborhood of is defined by a set (containing ) of all such that is a direct analytic continuation of satisfying where is a neighborhood of in That is,
In particular, a -neighborhood of (satisfying that ), is defined by the set (which contains ) of all such that is a direct analytic continuation of satisfying In other words, it is defined by A subset is called an open set, if either is an empty set or every point is an inner point, i.e., there is a -neighborhood of such that It is not hard to check that it defines a topology of
@@
Pick If then we can consider -neighborhoods of and of where Evidently, we have If for then we can take a function element which gives that in a smaller neighborhood of Using the uniqueness theorem of analytic functions, we obtain a contradiction with ∎
@@
The analytic continuation of elements in along a curve in is a continuous mapping of into namely, an analytic continuation of to along is a curve connecting and Moreover, if is an analytic continuation of along a curve, then is also an analytic continuation of along the same curve, i.e., the analytic continuation is symmetric.
@@
Evidently, It is sufficient to prove that Otherwise, we may assume that By Corollary 2.6, there exist exactly function elements at denoted by
Take small enough such that with Then, there is a unique integer such that Notice that the domain of definition of is for each one can redefine function element Combined with the definition of it yields immediately that can continue analytically to along i.e., is an analytic continuation of along However, it contradicts with the definition of Hence, we obtain The uniqueness of is immediate due to the uniqueness theorem of analytic functions. ∎
@@
Let be any two groups of distinct function elements at respectively. Then, can continue analytically to each other along any curve in
@@
Let be distinct function elements at every point With the help of Corollary 2.11, it is sufficient to prove that all can be continued analytically to each other. Let be the largest integer such that there are function elements in which continue analytically to each other along a curve. Note that Without loss of generality, one may assume that the function elements are just Then, it can be written as
where are holomorphic functions in a neighborhood of Using Vieta’s theorem, we obtain
It suffices to show that
can continue analytically to the whole
Apply Theorem 2.10, can continue analytically to along a curve for any where is a permutation of By Vieta’s theorem, determine the same in a small neighborhood of i.e., can continue analytically to Hence, holds. If then one can determine another set using the rest function elements However, it leads to that equation (7) is reducible, which is a contradiction. This completes the proof. ∎
@@
2.1.3. Properties of function elements near a multiple point
Let be a pair such that and
Let us take a small neighborhood of such that on Using the continuity of for there exists a number such that
@@ 2.1.3. Properties of function elements near a multiple point
According to Corollary 2.6, for any with there exist two groups of distinct meromorphic function elements at subordinate to equation (7), respectively. Apply Theorem 2.10, then for a curve with and we are able to extend analytically to along In other words, there is a permutation of such that can continue analytically to along respectively. The symmetry implies that can also continue analytically to along However, we should mention that won’t be, in general, continued analytically back to along respectively. In what follows, we would consider the analytic continuation of a function element along a closed Jordan curve surrounding any multiple point Take a smooth closed Jordan curve around such that which is close to and doesn’t surround or cross any connected components of that do not pass through Without loss of generality, we look at the analytic continuation of along starting from When returns to along it will turn to certain function If then one continues to extend analytically along in the same direction of rotation. Repeating this action, we shall see that cycles on and on periodically. Let be the smallest period such that can return to itself along all such curves Then, we get We call a branching point with order of for Repeatedly, it can be seen that is divided into finitely many groups. Let denote the number of groups. If the -th group has members, then
We just analyze the first group: Taking a suitable there exists a biholomorphic mapping such that
@@
where is the polydisc centered at 0 with polyradius in Setting where is the unit disc centered at 0 in the -plane. For any we have Set and write
We show that is a single-valued analytic function on Notice that is the branching order of we see that is holomorphic on because is so close to 0 that there exist no any branching points of in We only need to examine that is a removed singularity of To do so, it suffices to prove that is bounded near 0 due to Riemann’s analytic continuation. In fact, it follows from (7) and (9) that
@@
on If we obtain Thus, on i.e., is bounded on It means that is holomorphic on and it has the Taylor expansion
@@
at Let be the set of all branching points of in Say that a branching point of is non-singular, if does not lie in a crossing of components of and singular otherwise. Assume that is not a singular branching point of Then, is smooth if is small enough. Taking a suitable if necessary. Since is bounded on and defines the branching order of at which is independent of the choice of one can deduce from and Riemann’s analytic continuation that there exists a biholomorphic mapping with
@@
such that can be expanded locally into the following Pusieux series
at where are holomorphic functions in Comparing coefficients, we deduce that
@@
According to the above arguments, we show that is a -valued analytic function defined in a neighborhood of We call a -leaf function element. If one cuts along a real hypersurface passing through then can separate into single-valued analytic components, and all these components can turn to each other when they continue analytically along a closed curve around in that is sufficiently close to By the local expansion of we deduce that is a -multiple, -valued branching point with order of
Assume that is not a singular branching point of Set and define
By the previous arguments, one can also expand locally into a Pusieux series of the form
at If then can be expanded locally at in the form
We take a number with Set Then
It is clear that
Hence, the case turns to the case
@@
To conclude, for each there are distinct multi-leaf function elements subordinate to equation (7), denoted by where is a -valued meromorphic function with and Hence, has branching order at with When we say that is a branching point with order of Moreover, if is not a singular branching point of then we deduce that can expand locally into the convergent Pusieux series at That is, we have shown that
@@
Let be a -leaf function element subordinate to equation If is not a singular branching point of then can be expanded locally into the following convergent Pusieux series
at in a local holomorphic coordinate such that and for all where is the branching set of and are meromorphic functions.
Note that in Theorem 2.14 is a -multiple root of that is, can separate into distinct single-valued components with value at If then separates into distinct single-valued components which take at
According to the analytic continuation and uniqueness theorem, the above arguments show that equation (7) uniquely defines a -valued meromorphic function on outside the branching points of Let denote the set of all branching points of which is an analytic set in called the branching set of We generalize the notion of algebroid functions as follows.
@@
where are holomorphic functions locally defined on such that they define a meromorphic mapping
In a word, denote by the distinct single-valued meromorphic components of and given a -leaf function element subordinate to equation (7), with distinct single-valued components Then, there is a permutation of such that on
In the end, we provide an example to illustrate how to calculate the order of branching points of algebroid functions.
Example. Calculating the order of the -valued algebroid function
on at branching points
Since have single-valued components, respectively, we infer that has single-valued components; Since have single-valued components, respectively, we infer that has single-valued components. Thus, is a -valued algebroid function. Let stand for any closed Jordan curve in All the branching sets of are as follows
It is clear that
Order of
Take such that is surrounded by Note that are the only two branching sets of that pass through Therefore, is not allowed to surround or pass through any other branching sets of except for in order to determine the order of For all such we for
Thus, the order of is determined by the order of branching point of and the order of branching point of It is
Order of
Take such that is surrounded by Note that are the only two branching sets of that pass through Similarly, cannot surround or pass through any other branching sets of except for It implies that for
Thus, the order of is determined by the order of branching points of It is
Order of
Take such that is surrounded by Since are the only three branching sets of that pass through we have for
Thus, the order of is determined by the order of branching points of respectively. It is
Order of
Take such that is surrounded by Since are the only three branching sets of that passes through we have for
Thus, the order of is determined by the order of branching points of It is
2.2. Multi-leaf Complex Manifolds by Algebroid Functions
Let A neighborhood of is defined to be a set of all with such that is a direct analytic continuation of i.e.,
Take a suitable we have a biholomorphic mapping such that where is the unit ball centered at the coordinate origin 0 in Dfine a natural mapping by
Clearly, is a homeomorphism whose inverse is
Set
We prove that is a complex atlas of Let It suffices to examine that the transition mapping